# Resilient and Adaptive Replanning for Multi-Robot Target Tracking with Sensing and Communication Danger Zones

### v1

@@ Abstract
Multi-robot collaboration for target tracking presents significant challenges in hazardous environments, including addressing robot failures, dynamic priority changes, and other unpredictable factors. Moreover, these challenges are increased in adversarial settings if the environment is unknown. In this paper, we propose a resilient and adaptive framework for multi-robot, multi-target tracking in environments with unknown sensing and communication danger zones. The damages posed by these zones are temporary, allowing robots to track targets while accepting the risk of entering dangerous areas. We formulate the problem as an optimization with soft chance constraints, enabling real-time adjustments to robot behavior based on varying types of dangers and failures. An adaptive replanning strategy is introduced, featuring different triggers to improve group performance. This approach allows for dynamic prioritization of target tracking and risk aversion or resilience, depending on evolving resources and real-time conditions. To validate the effectiveness of the proposed method, we benchmark and evaluate it across multiple scenarios in simulation and conduct several real-world experiments.
@@ I Introduction
The multi-robot target tracking problem has been extensively studied for a wide range of applications, such as surveillance [1], environmental monitoring [2], and wildfire coverage [3]. Many works have focused on improving tracking accuracy, optimizing task allocation, and maximizing group efficiency within the constraints of limited resources[4, 5, 6, 7]. However, deploying these methods in real-world environments often fails due to the inherent dangers present in real-world scenarios. The tracked targets might adversarially attack robots, and the environment could contain hazardous factors, both of which can cause the failure of task-critical resources and affect the tracking performance of the robot team. To mitigate these effects, understanding risk and building resilience against failures are essential considerations in planning the actions of robots.
To address the risk of failures during target tracking, [8, 9, 10] developed a risk-aware tracking framework to balance team performance optimization with minimizing the risk of sensor failures. The authors in [11] extended the scenario by addressing the probabilistic communication failures. In this formulation, sensing danger zones interfere with the robots’ sensors and cause sensor failures, while communication danger zones disrupt normal communication between robots causing communication failures. However, these studies focus on risk awareness in known zones and lack resilient coordination strategies that enable recovery and re-coordination when team members suffer from failures.
Figure 1: Multi-robot target tracking with one communication danger zone (blue) and two sensing danger zones (red). (a) is the top view of the experiment, and (b) is a simulation with a similar setup. The red "+" represents a sensing attack, and the blue one represents a communication attack. The demo video is available at: https://youtu.be/ARMUzIKwsvc.
To fill these gaps, we propose a resilient and adaptive multi-robot target tracking framework for reactive task execution and recovery from failures that are caused by sensing and communication attacks. Our framework is resilient in the sense that robots can recover from sensing and/or communication failures to sustain tracking performance. The main contributions of this paper are as follows:
@@
  • We propose a resilient and adaptive strategy for replanning. Attacked robots that have recovered can share information about the danger zones with their teammates, which enables the remaining robots to adjust behaviors accordingly and reduce the likelihood of future failures.
  • We extensively evaluate our framework through both simulations and hardware experiments (as shown in Fig. 1), demonstrating its resilience and adaptability.
@@ II Related Works
Several approaches had been proposed for addressing multi-robot target tracking scenarios involving challenging hazards, which can be classified into two major categories. The first category focused on robustness against attacks [9, 8, 12, 13, 14, 15, 11, 10]. Robots adopted a stealthy motion planning paradigm to stay cautious beforehand and ensure that they can circumvent failures to an extent. For instance, [9] assumed the targets were hostile and proposed an approach to enable the robots to balance between optimizing tracking performance and ensuring safety. In [8], it decentralized the approach from [9] and designed a framework leveraging iterative consensus-based algorithms to track targets. The work in [12] designed a robust algorithm to secure robots against any number of sensing and communication attacks in the worst case. Another line of work investigated resiliency and explored how the robots could swiftly recover from failures, migrate attacks, continue task execution, and maintain satisfactory task performance [16, 17, 18, 19, 20, 21, 22]. In [17], the authors proposed an active information acquisition framework that can withstand any number of attacks and failures, even with minimal communication. Similarly,  [18] designed a scalable resilient multi-robot target tracking strategy with approximation bounds on tracking performance.
Our previous study in [23] focused on adaptation strategies to adjust the problem formulation. Likewise, the method in [9] adaptively balanced tracking performance with risk aversion, using the covariance matrix to encode tracking accuracy and the observability Gramian to measure risk. However, both approaches primarily addressed offline scenarios with pre-defined environmental information and were not suitable for real-time adaptation in dynamic settings. In dynamic cases, real-time adaptation of planning strategies was critical to improving effectiveness and success rates. Some planning frameworks considered safety by considering the clearance toward obstacles and dynamic limitations [24, 25]. In this work, we address this real-time adaptiveness by introducing optimization formulations with adaptive parameters.
@@ III-A Models of Dynamics and Danger Zones
We consider a team of robots, indexed as =$\mathcal{R}=\{1,2,\dots,M\}$, is tasked to track a team of targets indexed as $\mathcal{T}=\{1,2,\dots,N\}$. Each robot can estimate the target’s position relative to itself using a range and bearing sensor. The environment contains $p$ sensing danger zones $\{\mathcal{S}\}_{l=1}^{p}$ and $q$ communication danger zones $\{\mathcal{C}\}_{k=1}^{q}$. The location of each danger zone remains unknown and is only revealed if a robot encounters an attack in it. As soon as a robot enters a danger zone, it faces a certain probability of being attacked. The attacked robot can share information about the position of the danger zone across the communication network with its teammates. We also assume the attack effect is temporary, so the attacked robot recovers after it exits the danger zone. Dynamics models of robots and targets, as well as details of danger zones, are introduced below. Notably, in this paper, we use attack and failure interchangeably.
@@
(Robots and Targets Dynamics) Let $\mathbf{x}_{i,k}=[x_{i,t},~{}y_{i,t}]^{\top}$ denote the position of robot $i$, $i\in\mathcal{R}$ at time step $t$. The discrete-time dynamics of the robot is defined as:
@@
(Communication Danger Zone) The communication danger zone $\mathcal{C}_{k}$ is assumed to be static and defined with its center position $\mathbf{x}_{\mathcal{C}_{k}}$ that follows a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{\mathcal{C}_{k}},\mathbf{\Sigma}_{\mathcal{C}_{k}})$. The enemy in the danger zone can send noisy or deceiving signals to jam the robot such that its communication link becomes disabled. We use a similar definition as our prior work [11] based on distance ratio to define the communication danger zone, inside which robot $i$’s communication with its teammate $j$ at $\mathbf{x}_{j}$ is at risk of being jammed when the following condition holds:
@@
$\mathbb{P}(\mathbf{x}_{i}\in\mathcal{C}_{k})$ $=\int_{\lVert\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i}\rVert<\delta_{2}\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert}\mathrm{pdf}(\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i})d(\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i}).$ (6)
@@
III-B Chance-constrained Optimization before Failures
In our previous study [11], we assumed prior knowledge of the danger zones and focused on robustness against failures for target tracking. The tracking problem was formulated to ensure that all robots avoid danger zones while achieving the tracking tasks. We can enforce an upper bound on the probability of entering any danger zone in the form of chance constraints. Then, the multi-robot tracking problem at timestamp $t$ is formulated as:
@@ III-B Chance-constrained Optimization before Failures
where the objective function minimizes the weighted sum of tracking error and feasible control efforts with respective weights $w_{1}$ and $w_{2}$. Let the distance between robot $i$ to its farthest teammate be $c_{i}^{*}$, and the distance between robot $i$ and the jammer in the danger zone $\mathcal{C}_{k}$ be $a_{ik}$. The nonlinear optimization in Eq. 7 is approximated as:
@@
where $\mathbf{a}_{i,\mathcal{S}_{l}}=\mathbf{\mu}_{\mathcal{S}_{l}}-\mathbf{x}_{i}$, $\mathbf{a}_{i,\mathcal{C}_{k}}=\mathbf{\mu}_{\mathcal{C}_{k}}-\mathbf{x}_{i}$, and $\hat{\mathbf{a}}_{i,\mathcal{S}_{l}},~{}\hat{\mathbf{a}}_{i,\mathcal{C}_{k}}$ denote their corresponding normalized vectors. $\text{erf}(\cdot)$ is the standard error function. The function $f(\cdot)$ quantifies the error in target state estimation.
This problem formulation has a series of drawbacks: (i) It assumes that the prior locations of all danger zones are known beforehand; (ii) It adopts an over-conservative paradigm, in which robots never take risks for an improved target tracking quality; (iii) No systematic recovery procedure is provided for robots to retain decent target tracking performance once failures happen.
@@
IV Resilient and Adaptive Tracking
In our resilient target tracking framework, robots adopt different strategies to cope with sensing and communication attacks. Before encountering any attack from danger zones, the entire robot team initially has no prior knowledge of these zones and must be exposed to attacks to gather information. The robots use onboard range and bearing sensors to collectively estimate targets’ locations. The communication network transfers information instantaneously. With instantaneous information shared within the communication network, robots’ measurements are shared with all valid members, allowing the system to follow the optimal control strategy for maximizing tracking quality.
Given that the entire robot team is indexed as $\mathcal{R}$, we use $\mathcal{R}^{s}$ to denote indices for the subset of robots with normal sensing capabilities and $\mathcal{R}^{c}$ to represent indices for the subset of robots connected in the communication network. In other words, robots whose indices belong to $\mathcal{R}^{c}$ form a communication group. Initially, danger zones are unknown to all robots. We let $[p]^{d}$ denote indices for the set of sensing danger zones known to the robot performing individual planning, while $[p]^{d,c}$ represents indices for the set of sensing danger zones shared in the communication group. A similar definition applies to communication danger zones, where $[q]^{d}$ represents indices for the set known to the robot performing individual planning, and $[q]^{d,c}$ corresponds to the set shared within the communication group. For robots in the communication group, a centralized framework is proposed to share information and collectively optimize tracking performance. In contrast, when robots are under communication attacks or a combination of communication and sensing attacks, they need to perform individual planning and deploy different strategies.
@@ IV-A Danger Zone Attacks and Recovery Model
We assume that the damage caused by danger zones is temporary and can be removed under certain conditions. Robots can be attacked by both unknown and known danger zones. To effectively simulate the attack and recovery processes, we model a risk field associated with each danger zone to represent probabilistic attacks. Additionally, we define corresponding recovery conditions when robots are under sensing or communication attacks.
@@
Sensing recovery: In our problem setting, the attack is assumed to be temporary. If the probability of a robot being attacked is below a threshold $\mathbb{P}(\mathbf{x}\in\mathcal{S}_{l})\leq\epsilon_{l,0}$, we assume the robot will recover from sensing failure. The threshold $\epsilon_{l,0}$ may differ from $\epsilon_{1}$ (see Eq. 7c), allowing for varied recovery conditions and simulating model mismatches. For instance, a more aggressive recovery behavior can be achieved if we make $\epsilon_{l,0}$ to be smaller than $\epsilon_{1}$.
@@ IV-B Target Tracking without Communication Attack
The tracking error in the objective function is only for the robots with both sensing and communication abilities. $\nu_{i,l}$ and $\xi_{i,k}$ are slack variables.
IV-C Recovery and Adaptivness
@@ IV-C1 Sensing Attacks
Robots under sensing attacks are assumed to lose their sensing functionality but are still able to communicate with other robots. When a sensing attack occurs, the attacked robots determine the probabilistic distribution of the danger zone’s position, which is then broadcast to the rest of the team within the communication network instantly. Let $\mathcal{S}^{d}$ be the sensing danger zones that have just been discovered. Then, the set of known sensing danger zones revealed to the communication group is expanded, i.e. $[p]^{d,c}=[p]^{d,c}~{}\bigcup~{}[p]^{d}$, and constraints in Eq. 11 are adjusted accordingly.
@@ IV-C2 Communication Attacks
The robot under communication attack loses its communication capability and performs single-agent target tracking. It can track targets with its onboard measurements and update the estimate of targets’ positions. The robot can also trace back to estimate the origin of the attack. However, such an estimate cannot be broadcast to the rest of the team immediately since the attacked robot’s communication channel is disabled. Therefore, the task for a robot $r_{i}$ under communication attack would be planning for itself to maximize the tracking quality with its measurements and, at the same time, try to escape from the danger zone. After escaping from the danger zone, robot $r_{i}$ can broadcast the estimated location of the zone to the rest of the team.
Robots under communication failures estimate the positions of targets, optimize their trajectories to track targets, and independently navigate themselves out of the danger zone(s). Before the communication resumes, their teammates perform target tracking without knowing this communication danger zone. The optimization problem for each robot $r_{i}$, $i\in\mathcal{R}\setminus\mathcal{R}^{c}$ under communication attack is formulated as:

### v2

# Resilient Multi-Robot Target Tracking with Sensing and Communication Danger Zones

@@ Abstract
Multi-robot collaboration for target tracking in adversarial environments poses significant challenges, including system failures, dynamic priority shifts, and other unpredictable factors. These challenges become even more pronounced when the environment is unknown. In this paper, we propose a resilient coordination framework for multi-robot, multi-target tracking in environments with unknown sensing and communication danger zones. We consider scenarios where failures caused by these danger zones are probabilistic and temporary, allowing robots to escape from danger zones to minimize the risk of future failures. We formulate this problem as a nonlinear optimization with soft chance constraints, enabling real-time adjustments to robot behaviors based on varying types of dangers and failures. This approach dynamically balances target tracking performance and resilience, adapting to evolving sensing and communication conditions in real-time. To validate the effectiveness of the proposed method, we assess its performance across various tracking scenarios, benchmark it against methods without resilient adaptation and collaboration, and conduct several real-world experiments.
@@ I Introduction
The multi-robot target tracking problem has been extensively studied for a wide range of applications, such as surveillance [1], environmental monitoring [2], and wildfire coverage [3]. Most of the research has focused on improving tracking accuracy, optimizing task allocation, and maximizing group efficiency within the constraints of limited resources [4, 5, 6, 7, 8, 9]. However, deploying these methods in real-world environments often fails due to the inherent dangers present in real-world scenarios. The tracked targets might adversarially attack robots, and the environment could contain hazardous factors, both of which can cause the failure of task-critical resources and significantly degrade the tracking performance of the robot team. Mitigating these effects requires a thorough understanding of the associated risks, along with the development of resilient strategies against potential failures in planning the actions of robots.
To address the risk of failures during target tracking, [10, 11, 12] developed several risk-aware tracking frameworks to balance maximizing team performance with minimizing the risk of sensor failures. The authors in [13] extended the scenario by addressing the probabilistic communication failures. In these studies, sensing danger zones interfere with the robots’ sensors and cause sensor failures, while communication danger zones disrupt normal communication between robots, causing communication failures. However, these studies focus on risk awareness in known zones and lack resilient coordination strategies that enable recovery and re-coordination when team members suffer from failures.
Figure 1: Multi-robot target tracking with one communication danger zone (blue) and two sensing danger zones (red). (a) is the top view of the experiment, and (b) is a simulation with a similar setup. The red "+" represents a sensing attack, and the blue one represents a communication attack. The demo video is available at: https://youtu.be/0fq7fPuVaEA.
To bridge these gaps, we propose a resilient multi-robot target tracking framework for reactive task execution and recovery from failures that are caused by sensing and communication attacks. Our framework ensures resilience by enabling robots to recover from sensing and/or communication failures to sustain overall tracking performance. The main contributions of this paper are as follows:
@@
  • We propose a resilient coordination strategy. Attacked robots that have recovered can share information about the danger zones with their teammates, which enables the remaining robots to adjust their behaviors accordingly and reduce the likelihood of future failures.
  • We extensively evaluate our framework through both simulations and hardware experiments (Fig. 1), demonstrating its resilience and adaptability.
@@ II Related Works
In real-world scenarios, tracking multiple targets requires robot teams to be collectively resilient to external attacks and adaptive to evolving environmental conditions. Prior works addressing these challenges can be grouped into two categories. The first focuses on improving robustness against attacks [11, 10, 14, 15, 16, 17, 13, 12], often via stealthy motion planning to proactively circumvent failures. For example, [11] balances tracking and safety against hostile targets; [10] decentralizes this approach using consensus-based algorithms; [14] secures robots against any number of sensing and communication attacks.
The second category investigates resiliency, exploring how robots can quickly recover from failures, mitigate attacks, continue task execution, and maintain overall performance [18, 19, 20, 21, 22, 23, 24, 25]. For example, [19] proposes an active information acquisition framework that can withstand any number of attacks and failures, even with minimal communication. Similarly, [20] develops a scalable resilient tracking strategy with approximation guarantees, while [25] presents an information gathering control policy under unknown hazards, enabling adaptation as robots gradually explore and learn hazard locations.
To further enhance resiliency, our prior work [26] proposed adaptation strategies that modify the problem formulation as new danger zones are discovered. Similarly, [11] achieves a balance between tracking performance and risk aversion by leveraging the covariance matrix for tracking accuracy and the observability Gramian for measuring risk. However, these approaches primarily consider offline planning with pre-defined environmental information, making them less effective for real-time adaptation in unknown or changing settings. Since dynamic environments require continual adjustment of planning strategies [27, 28], our work introduces an optimization framework with soft chance constraints to enable flexible and resilient multi-robot behaviors in real time.
@@ III-A Models of Dynamics and Danger Zones
We consider a team of robots, indexed as =$\mathcal{R}=\{1,2,\dots,M\}$, is tasked to track a team of targets indexed as $\mathcal{T}=\{1,2,\dots,N\}$. Each robot can estimate the target’s position relative to itself using a range and bearing sensor. The environment contains $p$ sensing danger zones $\{\mathcal{S}\}_{l=1}^{p}$ and $q$ communication danger zones $\{\mathcal{C}\}_{k=1}^{q}$. The location of each danger zone remains unknown and is only revealed if a robot encounters an attack in it. As soon as a robot enters a danger zone, it faces a certain probability of being attacked. The attacked robot can share information about the position of the danger zone across the communication network with its teammates. We also assume the attack effect is temporary, so the attacked robot recovers after it exits the danger zone. Inter-robot collision is not considered because robots operate at different altitudes. Dynamics models of robots and targets, as well as details of danger zones, are introduced below. Notably, in this paper, we use attack and failure interchangeably.
@@
(Robots and Targets Dynamics) Let $\mathbf{x}_{i,k}=[x_{i,t},~y_{i,t}]^{\top}$ denote the position of robot $i$, $i\in\mathcal{R}$ at time step $t$. The discrete-time dynamics of the robot are defined as:
@@
(Communication Danger Zone) The communication danger zone $\mathcal{C}_{k}$ is assumed to be static and defined with its center position $\mathbf{x}_{\mathcal{C}_{k}}$ that follows a Gaussian distribution $\mathcal{N}(\mathbf{\mu}_{\mathcal{C}_{k}},\mathbf{\Sigma}_{\mathcal{C}_{k}})$. The enemy in the danger zone can send noisy or deceiving signals to jam the robot, disabling its communication link. We use a similar definition as our prior work [13] based on distance ratio to define the communication danger zone, inside which robot $i$’s communication with its teammate $j$ at $\mathbf{x}_{j}$ is at risk of being jammed when the following condition holds:
@@
$\mathbb{P}(\mathbf{x}_{i}\in\mathcal{C}_{k})$ $=\int_{\lVert\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i}\rVert<\delta_{2}\lVert\mathbf{x}_{i}-\mathbf{x}_{j}\rVert}\hskip-14.22636pt\mathrm{pdf}(\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i})d(\mathbf{x}_{\mathcal{C}_{k}}-\mathbf{x}_{i}).$ (6)
@@
We adopt Gaussian distributions to model the danger zones due to their mathematical tractability and their ability to naturally capture uncertainty about hazard locations. This continues, and probabilistic modeling enables analytical approximation of risk fields and facilitates real-time chance-constrained optimization.
III-B Chance-constrained Optimization before Attacks
In our previous study [13], we assumed prior knowledge of the danger zones and focused on robustness against potential attacks in target tracking. The tracking problem was formulated to ensure that all robots avoid danger zones while achieving the tracking tasks. We can enforce an upper bound on the probability of entering any danger zone as chance constraints. We focus on the optimization problem to find the next step control for the robot team for optimal tracking, following the setup in [8]. Then, the multi-robot tracking problem at timestamp $t$ is formulated as:
@@ III-B Chance-constrained Optimization before Attacks
where the objective function minimizes the weighted sum of tracking error in target state estimation $f(\cdot)$ and feasible control efforts with respective weights $w_{1}$ and $w_{2}$. The tracking error is a function of both robot states and the estimate of target positions, $\hat{\mathbf{z}}_{i,t+1}$, at time $t+1$. The $\epsilon_{1},\epsilon_{2}\in(0,1)$ are the confidence levels for sensing and communication failures, respectively. Let the distance between robot $i$ to its farthest teammate be $c_{i}^{*}$, and the distance between robot $i$ and the jammer in the danger zone $\mathcal{C}_{k}$ be $a_{ik}$. The nonlinear optimization in Eq. 7 is approximated as:
@@
[…]
This problem formulation has a series of drawbacks: (i) It assumes that the prior locations of all danger zones are known beforehand; (ii) It adopts an over-conservative paradigm, in which robots never take risks for improved target tracking quality; (iii) No systematic recovery procedure is provided for robots to retain target tracking performance under adversarial attacks.
@@
IV Resilient Target Tracking
In our resilient target tracking framework, robots adopt different strategies to cope with sensing and communication attacks. Before encountering any attack from danger zones, the entire robot team initially has no prior knowledge of these zones and must be exposed to attacks to gather information on these zones. The robots use onboard range and bearing sensors to estimate the targets’ locations collectively. The communication network transfers information instantaneously. With instantaneous information shared within the communication network, robots’ measurements are shared with all valid members, allowing the system to follow the optimal control strategy for maximizing tracking quality. We assume that the communication network operates without range constraints.
Given that the entire robot team is indexed as $\mathcal{R}$, we use $\mathcal{R}^{s}$ to denote indices for the subset of robots with normal sensing capabilities and $\mathcal{R}^{c}$ to represent indices for the subset of robots connected in the communication network. In other words, robots whose indices belong to $\mathcal{R}^{c}$ form a communication group. Initially, both types of danger zones are unknown to all robots. We let $[p]^{d}$ denote indices for the set of sensing danger zones known to the robot performing individual planning, while $[p]^{d,c}$ represents indices for the set of sensing danger zones shared in the communication group. A similar definition applies to communication danger zones, where $[q]^{d}$ represents indices for the set of communication danger zones known to the robot performing individual planning, and $[q]^{d,c}$ corresponds to the set of communication danger zones shared within the communication group.
For robots in the communication group, a centralized framework is proposed to share information and collectively optimize tracking performance. Robots that receive only sensing attacks lose their sensing abilities but can still stay in the communication group for information sharing. In contrast, when robots are under communication attacks or a combination of communication and sensing attacks, they need to perform individual planning and deploy different strategies.
@@ IV-A Danger Zone Attacks and Recovery Model
We assume that the damage caused by danger zones is temporary and can be recovered under certain conditions. Robots can be attacked by both unknown and known danger zones. To effectively simulate the attack and recovery processes, we model a risk field associated with each danger zone to represent probabilistic attacks. Additionally, we define corresponding recovery conditions when robots are under sensing or communication attacks.
@@
Sensing recovery: In our problem setting, the attack is assumed to be temporary. If the probability of a robot being attacked is below a threshold $\mathbb{P}(\mathbf{x}\in\mathcal{S}_{l})\leq\epsilon_{l,0}$, we assume the robot will recover from sensing failure. The threshold $\epsilon_{l,0}$ may differ from $\epsilon_{1}$ (see Eq. 7c), allowing for varied recovery conditions and simulating model mismatches. For instance, a more aggressive recovery behavior can be achieved if we make $\epsilon_{l,0}$ smaller than $\epsilon_{1}$.
@@ IV-B Target Tracking without Communication Attack
The tracking error in the objective function is only for the robots with both sensing and communication abilities. $\nu_{i,l}$ and $\xi_{i,k}$ are slack variables that relax the constraints, allowing robots to resiliently take risks and enter danger zones for better tracking qualities. Weights $w_{3}$ and $w_{4}$ are selected empirically to balance tracking performance and risk avoidance. Larger values result in more conservative behaviors, while smaller values allow riskier actions for improved tracking.
IV-C Resiliency and Recovery
@@ IV-C1 Sensing Attacks
Robots under sensing attacks are assumed to lose their sensing functionality but are still able to communicate with other robots. When a sensing attack occurs, the attacked robots determine the probabilistic distribution of the danger zone’s position, which is then broadcast to the rest of the team within the communication network instantly. Let $\mathcal{S}^{d}$ be the sensing danger zones that have just been discovered. Then, the set of known sensing danger zones revealed to the communication group is expanded, i.e. $[p]^{d,c}=[p]^{d,c}~\bigcup~\mathcal{S}^{d}$, and constraints in Eq. 11 are adjusted accordingly.
@@ IV-C2 Communication Attacks
The robot under communication attack loses its communication capability and performs single-robot target tracking and individual planning. It can track targets with its onboard measurements and update the estimate of the targets’ positions. The robot can also trace back to estimate the origin of the attack. However, such an estimate cannot be broadcast to the rest of the team immediately since the attacked robot’s communication channel is disabled. Therefore, the task for a robot $r_{i}$ under communication attack would be to plan to maximize the tracking quality with its measurements and, at the same time, try to escape from the danger zone. After escaping from the danger zone and regaining communication, robot $r_{i}$ can broadcast the estimated location of the zone to the rest of the team within the communication network.
Robots under communication attacks estimate the positions of targets, optimize their trajectories to track targets, and independently navigate themselves out of the danger zone(s). Before the communication resumes, their teammates perform target tracking without knowing this communication danger zone. The optimization problem for each robot $r_{i}$, $i\in\mathcal{R}\setminus\mathcal{R}^{c}$ under communication attack is formulated as: